# Belobog: Move Language Fuzzing Framework For Real-World Smart Contracts

### v1

@@ 3.4. Transaction Generation and Mutation
  • Step 1: Graph trace construction. The purpose is to walk on the type graph to find a graph trace $G_{trace}$, whose edge nodes are all default type nodes. Theoretically, we could start from any node in $G$ to find such a trace. But to make our algorithm run faster, we start from a function definition that only accepts void or primitive types as input. This choice does not affect the overall outputs of the algorithm. The reason is that, unless using shared objects from $\texttt{Pool}_{obj}$, the first function call in a valid transaction must create objects with void or primitive types. For example, the figure on the left in Figure 5 shows one graph trace obtained by walking on the example type graph in Figure 4(b) starting from loan<T>.
@@
  • Step 3: Transaction generation. With a concretized graph trace, we can then generate the final transaction by instantiating it. The process starts from all type nodes without incoming edges in the graph trace and adopts a width-first enumeration to generate each function call. For example, in Figure 5, we start from the type node u64 and generate the function call loan<u32>(1000), where 1000 is a random value generated according to its type. Then we generate the split_coin<u32>() and repay<u32>() function calls to consume all objects created by loan<u32>().
@@
  • Mutator 1: Value mutation. For primitive values in $S$, we mutate them by doing AFL-style havoc mutation(libafl), i.e., randomly mutating several bits or bytes in $S$.
@@
    If this search process fails, meaning that there is no valid type node in the original $G_{trace}$ or there is no such a path, we invoke the transaction generation algorithm with $f^{\prime}$ as the starting node to generate a transaction. This transaction will be appended to the original transaction to get a new transaction. For example, the dashed nodes in Figure 5 can also illustrate this process if we view swap<T1,T2> as the selected function node.
@@ 3.5. Concolic Execution
Concolic executor is an important component for ensuring Belobog reach critical paths in the target Move smart contracts. We implement the concolic executor on top of the MoveVM, which provides the necessary toolchains for us to obtain the runtime values of variables and stacks. In particular, stack values in MoveVM are also strongly typed, making it relatively easy for us to symbolically model vectors, which is notoriously hard for other languages such as C/C++ (symhard).
For a transaction, the general process of performing concolic execution on it consists of two steps: constraint collection and concolic mutation.
Constraint collection. We trace the execution of a transaction in its MoveVM bytecode and collect all the constraints along the execution path. For each function input that has a primitive type, Belobog attaches a symbolic value to it; For a vector of primitive type, Belobog attaches a vector of symbolic values to model its elements correspondingly. There are three major kinds of constraints that will be collected during execution, as detailed below.
  • Arithmetic operations. We collect arithmetic constrains on arithmetic operations, such as add and mul. In Move, such operations will abort if an overflow happens. Thus, Belobog will also include additional constraints to ensure that the solved values will never overflow.
  • Cast operations. Due to the type requirement, operands of an arithmetic operation need to be of the same type. MoveVM uses explicit cast bytecode, such as castu16, to change the operand type accordingly. At runtime, if a value exceeds the range of the target type, MoveVM will abort the execution. For example, if we cast a variable of type u64 to u32, it is abnormal if the runtime value of the variable exceeds the maximum value of u32. Belobog encode value constraints for cast operations by observing the source and destination types.
  • Vector access. When a transaction accesses a vector, the MoveVM will abort the execution if the access is out of the vector’s bounds. To avoid generating invalid values, Belobog encodes the index constraints for vector accesses.
Concolic mutation. Once Belobog collects all the constraints of a transaction, it will randomly pick one or more branches and flip their conditions in the constraints. The mutated constraints will be sent to an SMT solver. If a new set of inputs is found by the solver, the corresponding values in the original transaction will be replaced accordingly to obtain a new transaction. This new transaction will then be executed and added to the seed corpus if it increases code coverage.
We take the repay<T>() function in Figure 4(a) to illustrate how our concolic execution works and why it is necessary. Assume that the transaction is the one shown on the top right of Figure 5. The vulnerable path will only be explored if the assertion check paid==amount+fee in line  LABEL:line:fee is satisfied. Our symbolic executor will make the first argument of loan<T>() and the second argument of split_coin<T>() symbolic and hold the symbolic values SYM1 and SYM2, respectively. The symbolic executor will track the propagation of these symbolic values. Suppose that the paid=SYM2, amount = SYM1, and fee = 0.001*SYM1 in line LABEL:line:fee. To make the assertion pass, we need to satisfy the constraint SYM2==SYM1+0.001*SYM1. Our concolic executor can easily find the valid input values by assigning SYM2=1001 and SYM1=1000. However, the chance of getting valid input values would be negligible without concolic execution.
@@ 4. Implementation
We build Belobog from scratch, totaling 17k lines of Rust code. Specifically, we spend 4.1k lines to support forking a specific on-chain checkpoint to enable fast and local on-chain data access. With this support, Belobog can initialize the object pool with on-chain objects. The type graph building, graph tracing, and transaction generation and mutation take 3.7k lines, while the concolic mutation takes 2k lines.
Oracle. In order to detect bugs in Move smart contracts, we need oracles to decide if each transaction is executed correctly. Unfortunately, designing an automated and general oracle for fuzzing smart contracts has been a long-standing and unsolved problem (defiranger). The main reason is that oracles often depend on the business model or semantics of the target smart contract. One can think of it as detecting functional bugs in general software, which remains unsolved for decades. Therefore, following the similar practice in previous smart contract fuzzers like Foundry(foundry) and Echidna(echidna), Belobog provides complementary Move libraries to enable users to implement custom oracles in pure Move language by emitting specific events when invariants are violated. We will introduce more details in the evaluation in Section 5.
A recent empirical study (movescan) defines some common oracles for Move smart contracts. These oracles are mostly for checking mild issues in a smart contract, such as unnecessary type conversion, which may waste some gas. We implement all five oracles from (movescan) that can be used during fuzzing in Belobog. Additionally, we implement a new general oracle “earning profits” for detecting no-cost earnings. Below is a list of the details about these oracles.
@@
  • Precision Loss: check both operands of an arithmetic operation to determine if precision loss happens, e.g., 5/2 => 2 instead of 2.5.
@@
  • Unnecessary Bool Judgment: check if the operands of EQ or NEQ is a constant boolean value. In this case, the comparison is not necessary because the other operand itself could be used to determine the condition by either BrFalse or BrTrue.
  • SHL Overflow: detect left shift overflow.
  • Earning Profits: check if the amount of coins owned by the transaction sender increased after the execution. Since we do not add additional coins in generated transactions, the amount of coins should never be increased. Otherwise, we can earn profits without any cost.
Note that although we implement all these predefined oracles, we view them as the default oracles when users do not provide a better oracle. In our evaluation, we manually write oracles for the target smart contracts for a realistic experimental setup.
Supporting different blockchains. To date, the codebases and semantics of Sui and Aptos, two prevailing blockchains using Move, have diverged significantly. This leads to distinct features with entirely different interfaces. Thanks to the flexible and modular design of Belobog, we can easily support both blockchains with minor efforts. Thus, we believe our implementation can also be easily extended to other blockchains. For different blockchains, most changes are in coverage instrumentation and instruction tracing, totaling 483 and 638 lines of difference in the two blockchains, respectively.
@@ 5. Evaluation
  • RQ1 (Bug finding). Is Belobog effective in finding vulnerabilities in real-world Move smart contracts?
@@
  • RQ3 (Abalation study). How do the key components in Belobog affect the overall effectiveness?
  • RQ4 (Case study). Can Belobog reproduce real-world incidents?
@@
Dataset Construction. There are two datasets used in our evaluation. First dataset, DS1, is assembled from the audit reports provided by a well-known Web3 auditing company, BitsLab [^3] The company shared the anonymized data with us.. In this dataset, all Move smart contracts are audited manually by human experts with detailed vulnerability information. These reports provide us with ground-truth vulnerability locations to validate the bug-finding capability of fuzzers. Although it is possible that some vulnerabilities escape human experts’ audit, it does not affect the evaluation result of our experiments. Each vulnerability is labelled based on its severity. We select 37 projects that contain at least one Medium vulnerability in the past year. This yields 960 smart contracts with 9 Critical, 43 Major and 35 Medium vulnerabilities. [^4] Following MoveScan convention, we deem a Move module as a smart contract and one Move project usually contains multiple Move modules. Critical vulnerabilities mostly lead to the drain of liquidity or shares. Major vulnerabilities usually relate to the design flaws of the business models or improper access control that are not directly exploitable. Medium severity usually indicates a potentially incorrect setup or configuration causing minor losses or unexpected results, i.e., non-security issues. We exclude the Minor and Informational vulnerabilities because they are either code style issues or harmless defects, such as inconsistent events emitted or gas waste, and the audit reports mark these issues as not compulsory to fix. We attach the full anonymized contents of DS1 in our supplementary material. The second dataset, DS2, is from MoveScan (movescan). It contains 72 open-sourced Move projects, totalling 244 smart contracts. [^5] We exclude 1 incorrectly labeled project, 1 project that deletes its repository, and 3 projects that no longer build..
For DS1, since we have the ground-truth audit reports, we manually implement the oracle for each smart contract. Our oracle implementations are easy, averaging 15 lines per vulnerability. Since all projects in DS1 are deployed on Sui, our evaluation on DS1 only uses the Sui blockchain. For DS2, we enable all our predefined oracles and evaluate on both Sui and Aptos blockchains. Since the choice of blockchain does not affect the effectiveness of fuzzers, we do not explicitly separate them in our discussion.
@@
Baseline Fuzzers. We initially hoped to directly adopt two most relevant and state-of-the art smart contract fuzzers, ItyFuzz (ityfuzz) and Sui-Fuzzer (suifuzzer), as the baselines. Unfortunately, at the time of writing, there have been 6 months and 19 months since the last maintenance of ItyFuzz and Sui-Fuzzer, respectively. Our preliminary efforts showed that they failed to start their fuzzing process in most smart contracts. In order to evaluate them, we have to fix these two fuzzers before conducting the evaluation. Specifically, we add another 503 lines for ItyFuzz to upgrade its Sui dependency, resolve various dirty hacks that result in MoveVM invariant violations, and fix a few bugs causing segmentation faults and dead loops. For Sui-Fuzzer, we also upgrade its Sui dependency, fix its building process, and extend its support for the init function executed during Move contract deployment by another 381 lines of code.
@@
5.1. RQ1: Bug Finding
Table 1. Overall bug-finding results on DS1. Per-smart contract comparison between Belobog and ItyFuzz. ”MA” and ”ME” is abbreviated for ”Major” and ”Medium”, respectively.
@@ 5.1. RQ1: Bug Finding
Bug-finding results. Table 1 shows the bug-finding results of different fuzzers on the DS1. Belobog demonstrates the strongest bug-finding capability in all projects. Out of all 109 bugs, Belobog successfully detects more than 85% (93) of them. Notably, Belobog detects all the 9 Critical bugs, the most concerning vulnerabilities in the real world. On 24 out of 37 projects, Belobog finds all the bugs. In comparison, ItyFuzz only detects 3 bugs, with all of them being Major or Medium, and no Critical bugs can be found.
Regarding DS2, since it is created by a static analysis tool MoveScan, we only compare our tool with it. In total, Belobog finds 181 defects, while MoveScan reveals 176 defects with 20 false positives. The further breakdown of our predefined oracles and smart contracts on DS2 is available in our complementary appendix.
Sui-Fuzzer Deficiency. We do not include the result of Sui-Fuzzer in the table because it does not yield any findings even after we fix many bugs in it. Sui-Fuzzer always aborts early in our experiments because it does not support shared objects. We then investigate that even if Sui-Fuzzer could support shared objects, it would still fail all the fuzzing experiments because it still has several inherent design flaws: (1) it will forge objects of any types by mutating the objects unconditionally, leading to massive amount of false positives; (2) it only supports testing a single Move smart contract with selected functions while the real-world defects mostly span multiple ones; (3) it only supports filling up objects from the existing global owned objects, making it unable to correctly handle “Hot Potato” objects; and (4) it does not consider types of object and usually fill in objects with unmatched types.
Case study in DS1. We select one Major vulnerability, “Incorrect reserve calculation”, from the project 8bc50340 of DS1 to illustrate the design of our custom oracle and the effectiveness of our approach. [^6] The code and description are also anonymized. In general, a custom Move oracle of Belobog is designed to be a thin wrapper for each target function by logging relevant values and checking invariants. Figure 6 illustrates a simplified version of the custom oracle and original contract, where function oracle_buy<T>() is the manually written oracle for function buy<T>(). This smart contract provides a function buy<T>() to buy <T> with the <SUI> coins, a native currency provided by the standard library. The oracle function oracle_buy<T>() serves as a thin wrapper for buy<T>(), forwarding the function call while verifying that the reserve is appropriately decreased in line LABEL:line:reserve. When a transaction is to be executed, all of the function calls to buy<T>() will be replaced by oracle_buy<T>() to check oracle violations during the execution of this transaction. With this oracle, Belobog can detect this vulnerability in a few seconds once the input coin matches other constraints in the buy<T>() function. However, even with this oracle, ItyFuzz cannot generate a valid function call to buy<T>() due to not supporting the type parameters.
@@
2pub struct SUI has drop { ... } // struct representing the standard SUI coin
@@
4public fun buy<T>(payment: Coin<SUI>, amount: u64): (Coin<SUI>, Coin<T>) { ... }
@@
6public fun oracle_buy<T>(coin: Coin<SUI>, amount: u64): (Coin<SUI>, Coin<T>) {
7let previous_reserve = get_reserve();
8let out = buy(item_id); // Forward the parameters to the original implementation
@@
10report_oracle_violation("Incorrect reserve calculation.");
11}
12out
13}
Figure 6. An example oracle for detecting a Major vulnerability.
1
2fun count_leading_zeros(x: u16) {
3if (x & 0xFF00 == 0) { x = x << 8; };
4...
5}
Figure 7. A spurious left shift overflow MoveScan reports
False positives in MoveScan. MoveScan has 20 false positives while Belobog has no false positives. We take the Figure 7 as an example to illustrate how MoveScan produces false positives. Since MoveScan is a static analyzer, it does not know the actual value of each parameter and thus it will assume the maximum value for each variable to detect overflow. In this case, MoveScan assumes $x==65535$ and identifies $x=x<<8$ as a left shift overflow but ignores the guard x & 0xFF00 == 0, leading to a false positive. Belobog can always check the arithmetic at runtime and thus avoid any such false positives.
False negatives in Belobog. While Belobog shows strong capability of finding bugs on DS1, we still miss 15% (16) bugs for two reasons. First, some bugs are out of scope. For instance, some contracts are vulnerable to Denial-of-Service attacks because of iterating over too large vectors, while some other vulnerabilities relate to incorrect logs during execution. As a smart contract fuzzer, Belobog is not capable of detecting such vulnerabilities that do not affect the integrity of smart contracts. Second, some bugs need administrator privileges to trigger. For example, centralization risk is a common vulnerability where the administrator has too much power. Both Belobog and the baselines ItyFuzz and Sui-Fuzzer do not assume any special capabilities for testing. These bugs go beyond the capability of a dynamic fuzzer, and we consider them as one of our limitations. We will discuss more limitations in Section 6.
Figure 8. Belobog vs. ItyFuzz in terms of basic-block coverage on all projects from DS1. Each point represents one project. Belobog consistently achieves much higher coverage than ItyFuzz, indicating that it can explore deeper program states and reach more code.
Figure 9. Fuzzing throughput of Belobog and ItyFuzz on all projects from DS1. Throughput is measured in thousands of executed instructions per second (kI/s). Belobog executes significantly more instructions per second than ItyFuzz across all targets.

### v2

@@ 3.4. Transaction Generation and Mutation
  • Step 1: Graph trace construction. The purpose is to walk on the type graph to find a graph trace $G_{trace}$, whose source nodes (no incoming edges) and sink nodes (no outgoing edges) are all default type nodes. Theoretically, we could start from any node in $G$ to find such a trace. But to make our algorithm run faster, we start from a function definition that only accepts void or primitive types as input. This choice does not affect the overall outputs of the algorithm. The reason is that, unless using shared objects from $\texttt{Pool}_{obj}$, the first function call in a valid transaction must create objects with void or primitive types. For example, the figure on the left in Figure 5 shows one graph trace obtained by walking on the example type graph in Figure 4(b) starting from loan<T>.
@@
  • Step 3: Transaction generation. With a concretized graph trace, we can then generate the final transaction by instantiating it. The process starts from all type nodes without incoming edges in the graph trace and adopts a breadth-first enumeration to generate each function call. For example, in Figure 5, we start from the type node u64 and generate the function call loan<u32>(1000), where 1000 is a random value generated according to its type. Then we generate the split_coin<u32>() and repay<u32>() function calls to consume all objects created by loan<u32>().
@@
  • Mutator 1: Value mutation. For primitive values in $S$, we mutate them by doing AFL-style havoc mutation(Fioraldi et al., 2022), i.e., randomly mutating several bits or bytes in $S$.
@@
    If this search process fails, meaning that there is no valid type node in the original $G_{trace}$ or there is no such path, we invoke the transaction generation algorithm with $f^{\prime}$ as the starting node to generate a transaction. This transaction will be appended to the original transaction to get a new transaction. For example, the dashed nodes in Figure 5 can also illustrate this process if we view swap<T1,T2> as the selected function node.
@@ 3.5. Concolic Execution
Concolic executor is an important component for ensuring Belobog can reach critical paths in the target Move smart contracts. We implement the concolic executor on top of the MoveVM, which provides the necessary toolchains for us to obtain the runtime values of variables and stacks. In particular, stack values in MoveVM are also strongly typed, making it relatively easy for us to symbolically model vectors, which is notoriously hard for other languages such as C/C++ (Kapus and Cadar, 2019). We leave the implementation details in our supplementary materials.
@@ 4. Implementation
We build Belobog from scratch, totaling 17k lines of Rust code. Specifically, we spend 4.1k lines to support forking a specific on-chain checkpoint to enable fast and local on-chain data access. With this support, Belobog can initialize the object pool with on-chain objects. The type graph building, graph tracing, and transaction generation and mutation take 3.7k lines, while the concolic execution takes 2k lines.
Oracle. To detect bugs in Move smart contracts, we require oracles to determine whether each transaction is executed correctly. Unfortunately, designing an automated, general-purpose oracle for fuzzing smart contracts remains a long-standing open problem (Wu et al., 2023). The primary reason is that oracles often depend on the target smart contract’s business model or semantics. One can think of it as detecting functional bugs in general software, which remains unsolved for decades. Therefore, following a similar practice in previous smart contract fuzzers like Foundry(8) and Echidna(Grieco et al., 2020), Belobog provides complementary Move libraries to enable users to implement custom oracles in pure Move language by emitting specific events when invariants are violated. We will introduce more details in the evaluation in Section 5.
A recent empirical study (Song et al., 2024) defines some common oracles for Move smart contracts. These oracles are mostly for checking mild issues in a smart contract, such as unnecessary type conversion, which may waste some gas. We implement and refine all five oracles from (Song et al., 2024) that can be used during fuzzing in Belobog. Additionally, we implement a new general oracle “earning profits” for detecting arbitrary earnings. Below is a list of the details about these oracles.
@@
  • Precision Loss: check both operands of an arithmetic operation to determine if precision loss happens, e.g., 5/2 => 2 instead of 2.5. Belobog utilizes the concolic execution to precisely track the symbolic expressions rather than pattern matching (Song et al., 2024).
@@
  • Unnecessary Bool Judgment: check if at least one of the operands of EQ or NEQ is a constant boolean value. In this case, the comparison is not necessary because the other operand itself could be used to determine the condition by either BrFalse or BrTrue.
  • SHL Overflow: detect left shift overflow. Belobog tracks the runtime values instead of static analysis to ensure no false positives.
  • Earning Profits: check if the amount of coins owned by the transaction sender increased after the execution. Since we do not mint additional coins in generated transactions, the amount of coins should never be increased. Otherwise, we can earn profits without any cost, indicating serious issues.
Supporting different blockchains. To date, the codebases and semantics of Sui and Aptos, two prevailing blockchains using Move that once shared code, have diverged significantly. This leads to distinct features with entirely different interfaces. Thanks to the flexible and modular design of Belobog, we can easily support both blockchains with minor efforts. Thus, we believe our implementation can also be easily extended to other blockchains. For different blockchains, most changes are in coverage instrumentation and instruction tracing, totaling 483 and 638 lines of difference in the two blockchains, respectively.
@@ 5. Evaluation
  • RQ1 (Bug Reachability). Is Belobog effective at reaching vulnerabilities in real-world Move smart contracts?
@@
  • RQ3 (Ablation study). How do the key components in Belobog affect the overall effectiveness?
  • RQ4 (Incident Reproduction). Can Belobog reproduce previous on-chain incidents?
  • RQ5 (Real-world Application). Can Belobog find new bugs in real-world smart contracts?
@@
Dataset Construction. Two datasets are used in our evaluation. The first dataset, DS1, is assembled from audit reports provided by a well-known Web3 auditing company. In this dataset, all Move smart contracts are manually audited by human experts, with detailed vulnerability information and further confirmed by project developers. These reports provide us with ground-truth vulnerability locations to validate the bug-reachability capability of fuzzers. Although it is possible that some vulnerabilities escape human experts’ audit, it does not affect the evaluation results of our experiments. Each vulnerability is labeled based on its severity acknowledged by the developers. As the audit reports define, Critical vulnerabilities mostly lead to fund loss. Major vulnerabilities usually relate to the design flaws of the business models or improper access control that are not directly exploitable. Medium severity usually indicates a potentially incorrect setup or configuration causing minor losses or unexpected results, i.e., non-security issues. We exclude the Minor and Informational vulnerabilities because they are either code style issues or harmless defects, such as gas waste, and the audit reports mark these issues as not compulsory to fix. We select projects that contain at least one Medium vulnerability in the past year, which yields 37 projects containing 960 smart contracts with 9 Critical, 43 Major, and 35 Medium vulnerabilities. [^3] Following MoveScan convention, we deem a Move module as a smart contract and one Move project usually contains multiple Move modules. We attach the full anonymized DS1 in our supplementary materials. The second dataset, DS2, is from MoveScan (Song et al., 2024). It contains 72 open-source Move projects, totaling 387 smart contracts on both Sui and Aptos. [^4] We exclude 1 incorrectly labeled project, 1 project that deletes its repository, and 3 projects that no longer build..
As we discussed earlier in Section 4, designing an automated and general-purpose oracle for fuzzing smart contracts remains a long-standing open problem (Wu et al., 2023). Luckily, we have the ground-truth audit reports for DS1. We thus implement the oracle for each smart contract, averaging 15 lines per vulnerability. To avoid overfitting, we invite two additional Move auditors with at least 1 year of full-time experience to review the implementation. DS2 does not have ground truth about vulnerabilities, and we thus rely on our predefined oracles.
@@
Baseline Fuzzers. We initially hoped to directly adopt two most relevant and state-of-the-art Move smart contract fuzzers, ItyFuzz (Shou et al., 2023) and Sui-Fuzzer (27), as the baselines. Unfortunately, at the time of writing this paragraph, there have been 6 months and 19 months since the last maintenance of ItyFuzz and Sui-Fuzzer, respectively. Our preliminary efforts showed that they failed to start their fuzzing process in most smart contracts. To evaluate them, we must first fix these two fuzzers. Specifically, we add 503 lines to ItyFuzz to upgrade its Sui dependency, resolve various dirty hacks that cause MoveVM invariant violations, and fix a few bugs that cause segmentation faults and deadlocks. For Sui-Fuzzer, we also upgrade its Sui dependency, fix its build process, and extend its support for the init function executed during Move contract deployment by an additional 381 lines of code.
For DS1, where we use custom oracles, we evaluate all baseline fuzzers and analyzers with our constructed oracles on it. For DS2, where we can only use predefined oracles, since only MoveScan and Belobog support such oracles, we evaluate them on it. Additional evaluation breakdowns are provided in the supplementary materials.
@@
5.1. RQ1: Bug Reachability
Table 1. Overall bug-reachability results on DS1. Per-project comparison between Belobog and ItyFuzz. ”MA”, ”ME” and ”FP” are abbreviated for ”Major”, ”Medium” and ”False Positive”, respectively. Note Sui-Fuzzer is evaluated but excluded because it does not yield any true positive findings.
@@ 5.1. RQ1: Bug Reachability
Bug-reachability results. Table 1 shows the bug-reachability results of different fuzzers on the DS1. Belobog demonstrates the strongest bug-reachability capability in all projects. Out of all 109 bugs, Belobog successfully detects more than 85% of them. Notably, Belobog detects all the 9 Critical bugs, the most concerning vulnerabilities in the real world. On 23 out of 37 projects, Belobog finds all the bugs. In comparison, ItyFuzz only detects 3 bugs, with all of them being Major or Medium, and no Critical bugs can be found.
We further present the results of the predefined oracles against MoveScan on DS1 and DS2 in Table 2. The detailed breakdowns of all projects and oracles are available in our supplementary materials. Belobog outperforms MoveScan in both datasets and covers every finding MoveScan reports. On DS1 and DS2, Belobog covers additional 13 (11 Precision Loss, 1 Type Conversion, 1 Unnecessary Bool Judgement) and 28 (all Precision Loss) bugs, respectively. While MoveScan reports 8 and 12 false positive bugs on DS1 and DS2, Belobog has no false positive reports. We attribute such advantages to two reasons: (1) MoveScan scans these vulnerabilities within a single function context statically, while Belobog is capable of tracking values and invariants across multiple call boundaries at runtime; and (2) the concolic execution of Belobog can provide the precise context for the predefined oracles. Note that all the bugs in Table 1 cannot be detected by MoveScan because the oracles of MoveScan are only capable of detecting simple Move smart contract defects, while real-world vulnerabilities are usually complicated and bound to specific business models.
Sui-Fuzzer Deficiency. We do not include the result of Sui-Fuzzer in Table 1 because it does not yield any findings even after we fix many bugs in it. Sui-Fuzzer always aborts early in our experiments because it does not support shared objects. We then investigate that even if Sui-Fuzzer could support shared objects, it would still fail all the fuzzing experiments because it still has several inherent design flaws: (1) it will forge objects of any types by mutating the objects unconditionally, leading to massive amount of false positives; (2) it only supports testing a single Move smart contract with selected functions while the real-world defects mostly span multiple ones; (3) it only supports filling up objects from the existing global owned objects, making it unable to correctly handle “Hot Potato” objects; and (4) it does not consider types of objects and usually fills in objects with unmatched types.
@@
3public fun compute_boost_factor(vestar_amount: u128, farm_amount: u128, total: u128): u64 {
4// precision loss because division before multiplication
@@
6let divisor = mul_div(farm_amount, FACTOR * 2, total);
7((dividend / divisor) + FACTOR) / (FACTOR / 100)
8}
9public fun calculate_boost_weight(amount: u128, boost_factor: u64): u128 {
10// precision loss because division before multiplication
@@
12}
13public fun update_boost_factor(...) {
14let boost_factor = compute_boost_factor(...);
15let weight = calculate_boost_weight(asset_amount, boost_factor);
16}
Figure 6. A Precision Loss defect of the starswap-core-aptos project in DS2.
Table 2. Results of the predefined oracles on DS1 and DS2. The projects where Belobog and MoveScan report the same results are abbreviated. Full results are available in our supplementary materials.
DS1 Projects Belobog MoveScan DS2 Projects Belobog MoveScan
8bc50340 18 26 (8 FP) utils 0 6 (6 FP)
73c07115 1 0 originmate 39 16
e3564d47 29 27 pyth-crosschain-sui 2 8 (6 FP)
074ef533 2 0 Starswap-aptos farming 8 4
8f0c3e4c 1 0 LiquidswapRouterV2 1 0
fb8888cb 8 1
Total 816 811 (8 FP) Total 146 130 (12 FP)
Case Study in DS2. We demonstrate our advantage over MoveScan with two Precision Loss vulnerabilities from the same project from the starswap-core-aptos project of DS2 on the Aptos chain. The first Precision Loss happens at Line LABEL:line:loss1 in Figure 6. The root cause is that the division of total happens before the multiplication of FACTOR. The second Precision Loss is at Line LABEL:line:loss2 in Figure 6, where the division of FACTOR to boost_factor() happens in the function compute_boost_factor() before the multiplication of amount in calculate_boost_weight() when calling the function update_boost_factor(). MoveScan misses both vulnerabilities because it only detects within a single function call context and does not trace values across function boundaries. In addition, it only supports a very limited fixed pattern like (a / b) * c with exactly three operands. Belobog finds both Precision Loss vulnerabilities because we track symbolic expressions for each value across functions and accurately identify potential precision loss.
False negatives in Belobog. While Belobog shows strong capability of finding bugs on DS1, we still miss 15% (16) bugs for two reasons. First, some bugs are out of scope. For instance, some contracts are vulnerable to Denial-of-Service attacks due to iterating over too large vectors, while others are vulnerable to incorrect logging during execution. As a smart contract fuzzer, Belobog cannot detect vulnerabilities that do not affect the integrity of smart contracts. Second, some bugs need administrator privileges to trigger. For example, centralization risk is a common vulnerability where the administrator has too much power. Both Belobog and the baselines ItyFuzz and Sui-Fuzzer do not assume any special capabilities for testing. These bugs go beyond the capability of a dynamic fuzzer, and we consider them as one of our limitations. We will discuss more limitations in Section 6.
Figure 7. Edge coverage of Belobog and ItyFuzz on the projects from DS1.